OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions 7+√2x-1=10

OpenStudy (mertsj):

Subtract 7 from both sides first

OpenStudy (mertsj):

what do you get?

OpenStudy (mertsj):

\[\sqrt{2x-1}=3\]

OpenStudy (mertsj):

Now square both sides.

OpenStudy (mertsj):

\[2x-1=9\]

OpenStudy (mertsj):

\[2x=10, x=5\]